# Musical ethnocentrism in Large Language Models

### v1

@@ 2 Related work
Initial studies, such as those discussed in Tao et al. (2024) and Naous et al. (2023), show LLMs often encode biases favoring Western, English-speaking norms, impacting their fairness and representation of non-Western cultures as well as performance on non-Western topics, like Traditional Chinese Medicine Lingxuan Zhu and Luo (2024).
Further research seeks methodologies to measure and mitigate these biases more accurately. The interdisciplinary approach in Biedma et al. (2024) and the survey on modeling culture in LLMs Adilazuarda et al. (2024) propose new frameworks for understanding and adjusting the embedded cultural values in LLMs. The “CulturePark” Li et al. (2024a) initiative and the “NormAd” Rao et al. (2024) benchmark are notable in their attempts to simulate cross-cultural communication scenarios and assess LLMs’ adaptability to cultural contexts through synthetic story generation, providing novel approaches to evaluating cultural sensitivity and adaptability in AI technologies. The “CDEval” Wang et al. (2023) benchmark specifically addresses the need to evaluate the cultural dimensions of LLMs, integrating automated generation and human verification to assess cultural traits across multiple domains. Similarly, the “CultureLLM” culturellm2023 project aims to fine-tune LLMs on culturally diverse data.
In the music domain, Li et al. (2024b) illustrates domain-specific biases, arguing for more comprehensive benchmarks in varied knowledge domains.

### v2

@@ 2 Related work
Initial studies, such as those discussed in Tao et al. (2024) and Naous et al. (2023), show LLMs often encode biases favoring Western, English-speaking norms, impacting their fairness and representation of non-Western cultures as well as performance on non-Western topics, like Traditional Chinese Medicine Zhu et al. (2024).
Further research seeks methodologies to measure and mitigate these biases more accurately. The interdisciplinary approach in Biedma et al. (2024) and the survey on modeling culture in LLMs Adilazuarda et al. (2024) propose new frameworks for understanding and adjusting the embedded cultural values in LLMs. The “CulturePark” Li et al. (2024b) initiative and the “NormAd” Rao et al. (2024) benchmark are notable in their attempts to simulate cross-cultural communication scenarios and assess LLMs’ adaptability to cultural contexts through synthetic story generation, providing novel approaches to evaluating cultural sensitivity and adaptability in AI technologies. The “CDEval” Wang et al. (2023) benchmark specifically addresses the need to evaluate the cultural dimensions of LLMs, integrating automated generation and human verification to assess cultural traits across multiple domains. Similarly, the “CultureLLM” Li et al. (2024a) project aims to fine-tune LLMs on culturally diverse data.
In the music domain, Li et al. (2024c) illustrates domain-specific biases, arguing for more comprehensive benchmarks in varied knowledge domains.